A cyclist rides round a tract there times. Her friend uses a stop watch to record time at the start of the ride, after one circuit, and at the end of the there circuits. The readings from the stop watch are shown. 6,00 seconds at the start, 180,00 seconds after one circuit and 600,00 seconds at the end of three circuits. What is the average time for one circuit of the track?


1
Expert's answer
2021-03-31T09:38:59-0400

The time for the first circuit is


"t_1=180-6=174\text{ s}."

The time for the second and third laps is


"t_2+t_3=600-180=420\text{ s}."

The average is


"t=\frac13(t_1+t_2+t_3)=\frac13 (174+420)=198\text{ s}."